Report for March 2017

Report Date
March 2017

Report compiled by Tim Gray, IS Apps Programme Manager for Alex Carter  (Programme Owner)

Section One: Programme Commentary, RAG and Resources Update

Programme Commentary

In this reporting period work has continued on all active projects. An issue with the Bamboo Service has had an impact on two projects. A resource conflict has delayed completion of the Serengeti upgrade build, while the inability to deploy changes to MyEd channels has meant that COM032 Deployment Review is not signed off. However, on the latter we are looking to have this small deployment picked up by the MyEd Upgrade project.

Finally, the annual Wiki upgrade project has sped efficiently through planning, the brief will be signed off in early April. The project plans to start UAT in early May with a deployment to LIVE in  Mid-June.

RAG Status and Forward Look

Resource Summary

 Section Two: Projects Not Started

Project Info     Project Manager's Commentary
  All this year's projects are now in flight. 

 

Section Three: Projects in Progress

Project Info Project Manager's Commentary

COM031

Serengeti Document Managment Server Upgrade

Compliance

Milestones

07-Apr-2017

Build

14-Apr-2017

Integration

25-Apr-2017

Acceptance

26-Apr-2017

Upgrade Oracle CH

27-Apr-2017

Deploy to Live

05-May-2017

Deployment review

12-May-2017

Close

 

RAG

AMBER

Report by Sue Woodger for March 2017

Achievements

  • Making progress with the Dev issues – see Issues below

 Issues

  • Issues with the Dev build within this project continue with the redirect problem (piccl 4 describes in more detail). Following losing the DevTech resource to Priority 1 WEB009 we then lost him to the Bamboo issues further delaying the Build signoff and revised dates were submitted to WIS and subsequently approved.  This has had a knock effect on our other milestones but we are still confident that we can still meet closure date of 12 May as once the redirect is complete we need to get Netcall in to apply a patch to fix a security issue and they will also upgrade Oracle at the same time thereby once Build is approved we move everything into Test then after UAT into Live.

Next Steps

  • Build sign off by 7 April 2017

 

Project is Amber because we have to get Netcall in after we have tested the redirect issue to patch therefore not confident in 7 April date for Build signoff. Netcall will however upgrade to Oracle 12 at the same time.

COM032

Rollout Academic Timetable Integration with Office 365 Calendar

Discretionary

Milestones

Target Datesort descending Title
15-Feb-2017 Deployment Review - MyEd Portlet (with Opt-Out)
17-Apr-2017 Closure

RAG

AMBER

 

Report by Chris Copner for March 2017

Achievements in Last Period

  • Handover to Production Management complete

Issues

Issue

  • The deployment of the updated MyEd portlet and closure of this project  (JIRA COM032-21 and COM032-20) is dependent on resolution of the Bamboo issue (recorded as PICCL issue 3)

Next Steps

  • Deploy updated MyEd portlet
  • Close project

 

Project is amber due to issue 3 - cannot deploy updated MyEd portlet or close project until Bamboo issue resolved.

COM033 Personalised Email for Students and Alumni

Discretionary

Milestones

Target Datesort descending Title
31-Mar-2017 Planning Review
04-May-2017 Design Review
20-Jun-2017 Build Review
11-Jul-2017 Integration Review
08-Aug-2017 Acceptance Review
15-Aug-2017 Deliver
05-Sep-2017 Deployment Review
19-Sep-2017 Closure Review

RAG

GREEN

Report by Tim Gray March 2017

 

Achievements in Last Period

  • Planning activities completed
  • Planning Milestone achieved

Issues

  • Project budget not yet secured
  • No direct issues, however activities from build onwards have not yet been estimated

Next Steps 

  • Design workshops planned for early April

 

COM034 Annual Wiki Upgrade

Compliance

Milestones

07-Apr-2017 Brief sign off
28-Apr-2017 DEV build sign off
03-May-2017 Install to TEST
09-May-2017 Integration Sign off
09-Jun-2017 Acceptance Sign off
13-Jun-2017 Deploy to LIVE
27-Jun-2017 Deployment sign off review
14-Jul-2017 Project Closure

RAG Status

GREEN

 

 

Report by Chris Konczak for March 2017

 

Achievements in Last Period

  • We had the project kick off meeting to discuss the project brief
  • Brief written
  • Estimations completed
  • Extra budget of 13 days has been approved
  • Brief sign off meeting had where the brief was signed off clarification required on the database migration version
  • DEV environment fixed with data copied from LIVE.

Issues

  • None

Next Steps

  • Brief to be submitted to WIS
  • TEST environment data to be refreshed from LIVE
  • Install Wiki onto DEV environment and sign off the build

Section Four: Closed or Withdrawn Projects

Project Info Project Managers Commentary

COM026

Wiki Upgrade

Discretionary

 

Report by Tim Gray

This project was closed on 23/09/16  (WIS Notes 23/09/16)

Please see the Completion Report for further information

COM025

BI Architechture

Discretionary

Report by Craig Middlemass

This project was closed on 04/11/16  (WIS Notes 04/11/16)

Please see the Completion Report for further information

COM028

Upgrade onsite Exchange Mail Relay Servers

Discretionary

Report by Chris Copner

This project was closed on 11/11/16 (WIS Notes 11/11/16)

Please see the Completion Report for further information

COM022 Academic Timetable Integration into Office 365 Calendar  

Discretionary

Report by Chris Copner

This project was closed on 18/11/16  (WIS Notes 18/11/16)

Please see the Completion Report for further information

COM030 Document Management for IS HR

Discretionary

Report by Tim Gray

This project was closed on 27/01/17  (WIS Notes 27/01/17)

Please see the Completion Report for further information

AP67-005 COM BI Annual Service Pack Compliance

Withdrawn at Mid-Year Review

COM Mid-Year review 2016/17

Section Five: Actions/Notes from monthly meeting held on 10/04/2017

 

Project Action / Notes Responsibility Status
COM032 Tim and Chris Copner to WEB009 team to enquire how we can deploy independently of WEB009 schedule. Need to confirm that deployment plan is working. Agreement from Bill and Prod. 

Tim Gray

Chris Copner

In progress
SOA Replacement

Richard Good is leaving the University this week. Prior to his leaving a meeting should be held with DevTech to acsertain Richard's thoughts on SOA Decommisioning, planned for next year.

Mark Lang Open
COM033 Budget  

COM033 does not currently have adaquate budget allocated: 40d allocated v. 100d planning estimate. The estimate is likely to change after the Design. Tim to speak to Adam Wadee regarding days from other programmes. Alex also to investiagte how other options. 

Tim Gray

Alex Carter

Open - 
Dev Services Rep

Tim to follow up with Bill Lee on Dev Services rep at Programme meetings. Outcome: Due to scheduling clash Bill could not attend the April metting. He will attend nthe meeting in May.

Tim Gray Closed